Question 5

Without using finance calculator,

Noel bought a house by making a down payment of RM 45,000. The balance was borrowed from a bank which charged interest of 7% compounded monthly for 25 years with monthly payments of RM 1413.56.

(a) Find the amount borrowed.

(b) What is the market price of the house?

(c) If he failed to make the first 5 payments, how much should he pay on the 6th payment to settle all the outstanding arrears?
1
Expert's answer
2020-05-28T11:00:32-0400

a) "1,413.56=S*\frac{0.07*(1+0.07)^{25}}{(1+0.07)^{25}-1}"

"S=16,473.04"

b) "45,000+16,473.04=61,473.04"

c) "S=16,473.04*1.12^5=29,031.125"